# One Diffusion to Generate Them All

### v1

@@ Abstract
We introduce OneDiffusion, a versatile, large-scale diffusion model that seamlessly supports bidirectional image synthesis and understanding across diverse tasks. It enables conditional generation from inputs such as text, depth, pose, layout, and semantic maps, while also handling tasks like image deblurring, upscaling, and reverse processes such as depth estimation and segmentation. Additionally, OneDiffusion allows for multi-view generation, camera pose estimation, and instant personalization using sequential image inputs. Our model takes a straightforward yet effective approach by treating all tasks as frame sequences with varying noise scales during training, allowing any frame to act as a conditioning image at inference time. Our unified training framework removes the need for specialized architectures, supports scalable multi-task training, and adapts smoothly to any resolution, enhancing both generalization and scalability. Experimental results demonstrate competitive performance across tasks in both generation and prediction such as text-to-image, multiview generation, ID preservation, depth estimation and camera pose estimation despite relatively small training dataset. Our code and checkpoint are freely available at https://github.com/lehduong/OneDiffusion.
Figure 1: OneDiffusion is a unified diffusion model designed for both image synthesis and understanding across diverse tasks. It supports text-to-image generation (red box), conditional image generation from input images (orange box) and it’s reverse task Image understanding (orange box). It can also perform ID customization (blue box), and multi-view generation (purple box) with arbitrary number of input and output images.
@@ 1 Introduction
Diffusion models, particularly in text-to-image (T2I) generation, have recently shown remarkable results. Models such as DALL-E [43], Imagen [43], and Stable Diffusion [47, 41, 15] have established new benchmarks for generating high-quality, photorealistic images from text prompts. Additionally, recent studies have demonstrated the effectiveness of diffusion models in various other computer vision tasks, such as depth estimation [22] or optical flow estimation [49, 37], etc. However, despite these advancements, diffusion models are typically trained individually for either T2I generation or specific tasks.
In contrast, large language models (LLMs) (e.g. GPT-4 [1]) have demonstrated their ability to function as universal models. They can perform a wide range of tasks across different domains without the need for task-specific modules, and can effectively handle tasks they haven’t been explicitly trained on zero-shot. This universality has been immensely valuable; it has dramatically simplified using, training and scaling these models, and ultimately lead to better performance. This incentivizes us to ask whether diffusion models can become universal in a similar way.
Designing a unified architecture for diverse image synthesis tasks presents significant challenges. Current methods often depend on external add-ons to handle new tasks. For example, ControlNet [68] or T2I-Adapter [39] require a specialized module to encode the conditional inputs, and personalization models typically require encoding the identity through a pretrained facial recognition network and adding auxiliary losses to preserve identity [63, 58, 21]. Additionally, tasks vary widely in input requirements. For instance, multi-view generation alone requires handling arbitrary input-output view combinations, posed or unposed images, and camera pose conditioning [18, 50, 57, 34, 24], while image understanding tasks require diverse outputs such as depth, pose, or segmentation. Finally, existing training recipes are often tightly tuned to particular tasks and therefore cannot be relied on to generalize between tasks.
@@
To train our model we create the One-Gen dataset, which integrates high-quality data across a variety of sources. This includes standard T2I data along with synthetic outputs from state-of-the-art models to support a range of tasks such as depth estimation, segmentation, pose estimation, etc. Our dataset also incorporates data for ID customization and multiview generation, providing diverse conditioning setups. The One-Gen dataset enables scalable joint training across multiple tasks, flexible conditioning options, improved generalization, and serves as a solid foundation for our model.
To demonstrate how general-purpose our training algorithm is, we train OneDiffusion completely from scratch. First, we train on text-to-image to equip the model with general image synthesis abilities, then on One-Gen to learn the full set of tasks. Our final model has 2.8 billion parameters and is equipped with a diverse set of skill, shown in Figure 1. The model also adapts naturally to various resolutions, enabling zero-shot high-resolution generation for sequence tasks even when such resolutions were not encountered during training.
We evaluate OneDiffusion on a diverse set of tasks for both generative and predictive tasks. On T2I, OneDiffusion efficiently generates high-quality images while utilizing fewer number of parameters. In the multiview generation task, OneDiffusion demonstrates performance comparable to state-of-the-art methods that are specifically designed and exclusively trained for this purpose. On depth estimation, OneDiffusion exhibits performance that is competitive with SoTA diffusion-based depth estimation baselines. We also show OneDiffusion supports novel conditioning setups, such as text-to-multi-view and image-to-multi-view. For high-variability tasks like face identification from a single image, the model is capable of generating multiple consistent images featuring flexible expressions and poses, demonstrating strong generalization to unseen domains.
@@ 2 Related work
Diffusion models for generative task.
Recent advancements in diffusion models have greatly improved image generation capabilities, with models like Stable Diffusion [47, 41, 15, 4, 3, 72] setting new standards in text-to-image synthesis. Building on this foundation, controllable diffusion models such as ControlNet [68] and T2I-Adapter [39] extend functionality by enabling fine-grained control through auxiliary inputs like edge maps, depth maps, or human poses. Meanwhile, instruct-Pix2Pix [5] introduces natural language-guided image editing, making these tools more user-friendly. For personalized applications, identity-focused models, including IP-Adapter [63], InstantID [58], PhotoMaker [27], and PuLiD [21], personalize generation by conditioning on reference images. Meanwhile, in multi-view generation, recent methods [50, 57, 34, 18], employ camera ray embeddings or 3D geometry to achieve consistent viewpoints. Together, these innovations showcase the versatility of diffusion models in delivering controllable, personalized, and multi-perspective image synthesis.
Diffusion models for predictive tasks.
@@
Unified diffusion models.
@@ 3.2 Proposed Approach
Objective.
We cast the problem of image generation with multimodal conditions as sequential modeling. Inspired by previous work on diffusion model for sequential data [69, 48, 7], we jointly model all conditions and target images as a sequence of “views”. Note that the number of views $N$ is determined by tasks. Particularly, $N=1$ for text-to-image tasks, $N=2$ for image-to-image translation such depth/pose/image editing etc, $N>2$ for multiview generation or ID customization.
@@ 3.3 Implementation Details
Text-to-Image (1 views).
@@
Image-to-Image (2 views).
@@
Multiview Generation (4-12 views).
@@ 4 One-Gen Datasets
For simpler tasks like deblurring, inpainting, image generation from canny edge, or upscaling, we use a $1M$-sample subset of our synthetic data and apply the relevant pre-processor for each image to create an input condition for it. For more complex tasks, we create a synthetic dataset from outputs generated by Midjourney, Stable Diffusion [15], and Flux-dev following the process outlined below:
@@ 5 Experiments
In this section, we evaluate our OneDiffusion model on broad range of image generation and understanding tasks. We do not perform task-specific finetuning in any results. Details about additional qualitive exampels are in Appendix.
@@ 5.2 Controllable Image generation
We show the experiment with image-to-image translation using various source domains, including HED, depth map, human pose, semantic map, bounding boxes. We report the qualitative results in Figure 4 and Figure 19 in appendix. The results demonstrate that OneDiffusion effectively aligns with the conditioning image across a wide range of input conditions during the generation process, leveraging pure attention mechanisms and supplementary information from captions.
@@ 5.3 Multiview Generation
We evaluate the multiview generation capabilities of our method using the Google Scanned Object dataset [13]. Table 2 presents a comparison of our approach (OneDiffusion) with state-of-the-art methods, including Zero123 [32], Zero123-XL [12], and EscherNet [15]. It is important to note that these models are specifically designed and exclusively trained for the multiview generation task. In contrast, our model supports a variable number of conditional inputs and, due to its flexible denoising framework, can incorporate additional conditional views even when their camera poses are unknown.
As shown in Table 2, OneDiffusion consistently outperforms Zero123 and Zero123-XL in the 1-view condition. Additionally, OneDiffusion demonstrates strong performance even when camera poses are unknown. For instance, in the 2-view condition with unknown poses, OneDiffusion achieves a PSNR of 19.83, which is only marginally lower than the PSNR of 20.22 obtained when camera poses are known. Similarly, in the 3-view condition, OneDiffusion achieves PSNR values of 20.64 and 21.79 for unknown and known poses, respectively. These results highlight the adaptability and effectiveness of our method across varying input conditions, demonstrating its potential for practical applications in multiview generation.
Figure 5 presents two qualitative examples of our multiview generation from a single front-view image. Our model successfully generates consistent views across varying azimuths and elevations. For additional visualizations, please refer to Figures 10 and 11 in Appendix. As a result of the flexibility of our framework, we can also perform text-to-multiview directly by “masking” all images and only inputing the camera poses as in Figure 12 in Appendix.
@@ 5.5 Depth Estimation
Figure 7: Qualitative comparison between a diffusion-based depth estimation - Marigold [22] and our methods.
@@
Table 4: Comparison of depth estimation methods on NYUv2 and DIODE datasets.
For image understanding tasks, we evaluate our model’s performance on monocular depth estimation using standard benchmarks: NYUv2 [51] and DIODE [56]. We report the quantitative results in Table 4.Quantitative results are presented in Table 4, where our model achieves competitive performance compared to baselines that leverage pretrained text-to-image diffusion models, such as Marigold [22]. However, as shown in Figure 7, our model demonstrates greater robustness than diffusion-based depth estimators like Marigold. Specifically, it excels in handling open-world images, including paintings, hazy weather, and unconventional textures. For further qualitative comparisons, refer to Figures 15 and 16.
@@ 5.6 Camera Pose Estimation
5.7 Other Tasks
Since directly extracting masks, bounding boxes, and keypoints from raw output images is not straightforward, we provide several qualitative results for human pose estimation and semantic segmentation on COCO dataset in Figures 17 and 18 in the appendix, respectively. Since our model does not distinguish tasks of conditions and images during training, its performance on understanding tasks serves as an additional indicator of the model. We leave further exploration of this aspect for future work.
6 Conclusion
Supplementary Material

### v2

@@ Abstract
We introduce OneDiffusion, a versatile, large-scale diffusion model that seamlessly supports bidirectional image synthesis and understanding across diverse tasks. It enables conditional generation from inputs such as text, depth, pose, layout, and semantic maps, while also handling tasks like image deblurring, upscaling, and reverse processes such as depth estimation and segmentation. Additionally, OneDiffusion allows for multi-view generation, camera pose estimation, and instant personalization using sequential image inputs. Our model takes a straightforward yet effective approach by treating all tasks as frame sequences with varying noise scales during training, allowing any frame to act as a conditioning image at inference time. Our unified training framework removes the need for specialized architectures, supports scalable multi-task training, and adapts smoothly to any resolution, enhancing both generalization and scalability. Experimental results demonstrate competitive performance across tasks in both generation and prediction such as text-to-image, multiview generation, ID preservation, depth estimation and camera pose estimation despite a relatively small training dataset. Our code and checkpoint are freely available at https://github.com/lehduong/OneDiffusion.
Figure 1: OneDiffusion is a unified diffusion model designed for both image synthesis and understanding across diverse tasks. It supports text-to-image generation (red box), conditional image generation from input images (orange box) and it’s reverse task Image understanding (green box). It can also perform ID customization (blue box), and multi-view generation (purple box) with arbitrary number of input and output images.
@@ 1 Introduction
Diffusion models, particularly in text-to-image (T2I) generation, have recently achieved remarkable results. Models such as DALL-E [46], Imagen [46], and Stable Diffusion [50, 44, 15] have established new benchmarks for generating high-quality, photorealistic images from text prompts. Additionally, recent studies have demonstrated the effectiveness of diffusion models in various other computer vision tasks, such as depth estimation [23] or optical flow estimation [53, 38], etc. However, despite these advancements, diffusion models are typically trained individually for either T2I generation or specific tasks.
In contrast, large language models (LLMs) (e.g. GPT-4 [1]) have demonstrated their ability to function as universal models. They can perform a wide range of tasks across different domains without the need for task-specific modules, and can effectively handle tasks they have not been explicitly trained in a zero-shot manner. This universality has been immensely valuable; it has dramatically simplified using training and scaling these models, and ultimately led to better performance. This incentivizes us to ask whether diffusion models can become universal in a similar way.
Designing a unified architecture for diverse image synthesis tasks presents significant challenges. Current methods often depend on external add-ons to handle new tasks. For example, ControlNet [73] or T2I-Adapter [40] require specialized modules to encode the conditional inputs, and personalization models typically require encoding the identity through a pretrained facial recognition network and adding auxiliary losses to preserve identity [68, 63, 21]. Additionally, tasks vary widely in their input requirements. For instance, multi-view generation alone requires handling arbitrary input-output view combinations, posed or unposed images, and camera pose conditioning [18, 54, 62, 35, 25], while image understanding tasks require diverse outputs such as depth, pose, or segmentation. Finally, existing training recipes are often tightly tuned to particular tasks and therefore cannot be relied on to generalize between tasks.
@@
To demonstrate how general purpose our training algorithm is, we train OneDiffusion completely from scratch. First, we train on text-to-image task to equip the model with general image synthesis abilities, then on our One-Gen dataset to learn the full set of tasks. Our final model has 2.8 billion parameters and is equipped with a diverse set of skills, shown in Figure 1. The model also adapts naturally to various resolutions, enabling zero-shot high-resolution generation even when such resolutions were not encountered during training.
We evaluate OneDiffusion on a diverse set of both generative and predictive tasks. On T2I, OneDiffusion efficiently generates high-quality images while utilizing fewer number of parameters. In the multiview generation task, OneDiffusion demonstrates performance comparable to state-of-the-art methods that are specifically designed and exclusively trained for this purpose. We also show that OneDiffusion supports novel conditioning setups, such as text-to-multi-view and image-to-multi-view. For high-variability tasks like face identification from a single image, the model is capable of generating multiple consistent images featuring diverse expressions and poses, demonstrating strong generalization to unseen domains.
@@ 2 Related work
Diffusion models for generative tasks
Recent advancements in diffusion models have greatly improved image generation capabilities, with models like Stable Diffusion [50, 44, 15, 4, 3, 77] setting new standards in text-to-image synthesis. Beyond general image generation, controllable diffusion models such as ControlNet [73] and T2I-Adapter [40] allow fine-grained control via auxiliary inputs like edge or depth maps. Similar structured conditioning has been applied to inverse problems [57, 42, 43], enabling applications such as super-resolution or inpainting. Meanwhile, instruct-Pix2Pix [5] introduces natural language-guided image editing, making these tools more user-friendly. For personalized applications, identity-focused models, including IP-Adapter [68], InstantID [63], PhotoMaker [28], and PuLiD [21], personalize generation by conditioning on reference images. Moreover, in multi-view generation, recent methods [54, 62, 35, 18], employ camera ray embeddings or 3D geometry to achieve consistent viewpoints. Together, these innovations showcase the versatility of diffusion models in delivering controllable, personalized, and multi-perspective image synthesis.
Diffusion models for predictive tasks
@@
Unified diffusion models
@@ 3.2 Proposed Approach
Objective
We cast the problem of image generation with multimodal conditions as sequential modeling. Inspired by previous work on diffusion model for sequential data [74, 51, 7], we jointly model all conditions and target images as a sequence of “views”. Note that the number of views $N$ is determined by tasks. Particularly, $N=1$ for text-to-image tasks, $N=2$ for image-to-image translation such depth/pose/image editing, etc, $N>2$ for multiview generation or ID customization.
@@ 3.3 Implementation Details
Text-to-Image (1 views)
@@
Image-to-Image (2 views)
@@
Multiview Generation (4-12 views)
@@ 4 One-Gen Datasets
For simpler tasks e.g. deblurring, inpainting, image generation from canny edge, or upscaling, we use a $1M$-sample subset of our synthetic data and apply the corresponding pre-processor for each image to create an input condition. For more complex tasks, we create a synthetic dataset from outputs generated by Midjourney, Stable Diffusion, and Flux-dev following the below process:
@@ 5 Experiments
We evaluate our OneDiffusion model on broad range of image generation and understanding tasks. We do not perform task-specific finetuning in any results. Details about additional qualitive examples are in the Appendix.
@@ 5.2 Controllable Image generation
We show the experiment with image-to-image translation using various source domains, including HED, depth map, human pose, semantic map, bounding boxes. We report the qualitative results in Figure 4 and 19 in appendix. Generated images of OneDiffusion consistently conform various types of conditions by purely utilizing attention mechanisms and supplementary information from captions.
@@ 5.3 Multiview Generation
We assess our method’s multiview generation capabilities using the Google Scanned Object dataset. Table 2 compares our approach (OneDiffusion) with state-of-the-art methods like Zero123, Zero123-XL, and EscherNet, which are tailored for multiview tasks. Unlike these, OneDiffusion supports variable conditional inputs and can handle additional views with unknown camera poses due to its flexible denoising framework.
In Table 2, OneDiffusion outperforms Zero123 and Zero123-XL in the 1-view condition and maintains strong results with unknown poses, e.g., a PSNR of 19.83 (2-view, unknown) vs. 20.22 (known), and 20.64 (3-view, unknown) vs. 21.79 (known). Figure 5 shows consistent multiview outputs from a single front-view image, with more examples in Appendix Figures 10 and 11. Our framework also enables text-to-multiview generation using only camera poses, as shown in Figure 12.
@@ 5.5 Depth Estimation
Table 4: Comparison of depth estimation methods on NYUv2 and DIODE datasets. OneDiffusion achieves competitive performance compared to previous depth estimation methods.
For image understanding tasks, we evaluate our model’s performance on monocular depth estimation using standard benchmarks: NYUv2 [55] and DIODE [61]. We report the results in Table 4. Our model achieves competitive performance compared to baselines that leverage pretrained text-to-image diffusion models, such as Marigold [23]. Notably, as illustrated in Figures 15 and 16, our model demonstrates superior robustness than diffusion-based depth estimators like Marigold. Specifically, it excels in handling open-world images, including paintings, hazy weather, and unconventional textures.
6 Conclusion
7 Acknowledgements
Stephan Mandt acknowledges support from the National Science Foundation (NSF) under an NSF CAREER Award IIS-2047418 and IIS-2007719, the NSF LEAP Center, by the Department of Energy under grant DE-SC0022331, the IARPA WRIVA program, the Hasso Plattner Research Center at UCI, the Chan Zuckerberg Initiative, and gifts from Qualcomm and Disney.
Supplementary Material
8 Additional quantitative results
@@ 8.1 Camera Pose Estimation
8.2 Image Editing and Subject-driven generation
We evaluate the performance of OneDiffusion for instruction-based image editing with the PIE-Bench dataset [22] in Table 6; and for subject-driven generation using DreamBench [52] in Table 7. OneDiffusion achieves strong performance compared to specialized editing and generation approaches without any fine-tuning.
Method Background Preservation CLIP Semantics
PSNR $\uparrow$ LPIPS $\downarrow$ Whole $\uparrow$ Edited $\uparrow$
Prompt-to-Prompt 17.87 208.80 25.01 22.44
Null-text Inversion 27.03 60.67 24.75 21.86
PnPInversion 27.22 54.55 25.02 22.10
Pix2pix-zero 20.44 172.22 22.80 20.54
MasaCtrl 22.17 106.62 23.96 21.16
InstructPix2Pix 16.69 271.33 23.49 22.20
MGIE 21.20 142.25 24.28 21.79
EditAR 21.32 117.15 24.87 21.87
OneDiffusion 27.49 56.67 25.84 22.34
Table 6: Evaluating image editing on PIE-Bench.
Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
Real Images (Oracle) 0.774 0.885
Fine-Tuning
Textual Inversion 0.569 0.780 0.255
DreamBooth 0.668 0.803 0.305
BLIP-Diffusion 0.670 0.805 0.302
Tuning Free (Zero-shot)
Re-Imagen 0.600 0.740 0.270
BLIP-Diffusion 0.594 0.779 0.300
OneDiffusion 0.692 0.814 0.297
Table 7: Evaluating subject-driven generation on DreamBench.
@@ 9 Additional qualitative results
Zero-shot task composition
OneDiffusion demonstrates remarkable generalization capabilities during inference, extending beyond its training on single-condition images to handle multiple conditioning inputs. Notably, it performs zero-shot task composition, such as inpainting with a reference face or generating images based on semantic maps and human pose, as illustrated in Figure 8.
Figure 8: OneDiffusion is capable of performing several zero-shot task compositions. Left is inpainting with reference image and right is image generation with human pose and semantic map (butterfly, flower, stone)
10 Summary Datasets